OpenStudy (ttp):

Which literal equation is equivalent to b equals five a minus 4? a equals four b plus five a equals minus five plus four b a equals b plus four divided by five a equals b minus four divided by five

OpenStudy (anonymous):

a equals b plus four divided by five becouse given is b=5-4a then a=(b+4)/5

OpenStudy (ttp):

Thanks!

OpenStudy (ttp):

:)

OpenStudy (anonymous):

Its ok....@TTP